OpenStudy (anonymous):

find an equation of the line that passes through (3,5)and is perpendicular to y=3x-7

OpenStudy (aaliyahfairgood):

slope int. form of equation y-y1=m(x-x1) y-5=-1/3(x-3) -1/3 because perpendicular slopes are negative reciprocals y-5=-1/3x+1 add 5 to both sides for equality y=-1/3x+6